ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

8. Analysis

1991 MS Classification: Primary 47D06; Secondary 47A50 Lizama, Carlos, Departamento de Matemática, Universidad de Santiago de Chile A Characterization of Uniform Continuity for Volterra Equations in Hilbert Spaces We show that the norm continuity of the resolvent for a Volterra equation of scalar type is equivalent to the decay to zero of an holomorphic operator family along some imaginary axis. The result extends a characterization of eventually norm continuous semigroups obtained by P. You (Proc. Amer. Math. Soc. 116 (1992)).